OpenStudy (anonymous):

Which of the following is the solution to the equation sqrt5^8n = 125^(n +4) n = 12 n = 4 n = –4 n = –12

OpenStudy (anonymous):

sqrt5^8n=5^4n 125^(n +4)=5^3(n+4) 5^3(n+4)=5^4n 3(n+4)=4n

OpenStudy (anonymous):

shfanks :)